# SAINT: Attention-Based Modeling of Sub-Action Dependencies in Multi-Action Policies

### v1

@@ Abstract
The combinatorial structure of many real-world action spaces leads to exponential growth in the number of possible actions, limiting the effectiveness of conventional reinforcement learning algorithms. Recent approaches for combinatorial action spaces impose factorized or sequential structures over sub-actions, failing to capture complex joint behavior. We introduce the Sub-Action Interaction Network using Transformers (SAINT), a novel policy architecture that represents multi-component actions as unordered sets and models their dependencies via self-attention conditioned on the global state. SAINT is permutation-invariant, sample-efficient, and compatible with standard policy optimization algorithms. In 15 distinct combinatorial environments across three task domains, including environments with nearly 17 million joint actions, SAINT consistently outperforms strong baselines. [^1] Code is available at https://github.com/matthewlanders/SAINT
@@ 1 Introduction
Reinforcement learning (RL) has achieved remarkable success across a range of domains, primarily through methods designed for either small, discrete action spaces [15, 28, 41, 27], such as those in classic games, or continuous control [11, 12, 22, 33] typical of robotics. However, many real-world problems fall between these extremes, involving large, discrete combinatorial action spaces defined as Cartesian products of multiple subspaces, with each joint action comprising several coordinated sub-actions: $\mathbf{a}=(a_{1},\dots,a_{A})$. Such settings arise in critical applications including traffic signal control [32], drug selection in healthcare [35], and job scheduling in multi-resource compute clusters [24].
Traditional RL methods typically model discrete action spaces using a flat categorical policy. In combinatorial settings, this becomes intractable as the number of possible actions grows exponentially with the number of sub-actions, scaling as $\prod_{d=1}^{A}m_{d}$ for $A$ sub-action dimensions, where $m_{d}$ denotes the number of discrete choices in the $d$-th dimension. To mitigate this combinatorial explosion, existing approaches [8, 31, 38, 47] often rely on simplifying assumptions that constrain the representational capacity of the policy class. A common strategy assumes conditional independence among sub-actions given the state, factorizing the policy as $\pi(\mathbf{a}|s)=\prod_{i}\pi_{i}(a_{i}|s)$ [31, 38]. While this provides computational efficiency, it fails to capture cooperative or antagonistic interactions between sub-actions, limiting the policy’s ability to express complex joint behaviors. An alternative is to impose a fixed autoregressive order, modeling the policy as $\pi(\mathbf{a}|s)=\prod_{i}\pi_{i}(a_{i}|s,a_{<i})$ [47]. Although more expressive, this approach introduces an arbitrary sequence over sub-actions — even in domains where such orderings are not meaningful — breaking permutation invariance and potentially impairing learning when the imposed order misaligns with the true dependency structure.
Many real-world tasks violate both the independence and fixed-order assumptions. For example, in healthcare, drug combinations often exhibit complex interaction effects: two drugs may be beneficial individually but harmful together. Accurately modeling such dependencies requires policy architectures that can capture joint interactions without relying on conditional independence or arbitrary sequencing assumptions. This motivates the development of more flexible, permutation-invariant models for combinatorial action spaces.
We propose the Sub-Action Interaction Network using Transformers (SAINT), a novel policy architecture that models multi-component actions as unordered sets of sub-actions. SAINT combines principles from permutation-invariant set processing with the Transformer’s self-attention mechanism [21, 43] to capture high-order dependencies among sub-actions. To incorporate global state information, SAINT first applies Feature-wise Linear Modulation (FiLM) [30] to the initial sub-action embeddings. These state-conditioned embeddings are then processed by Transformer blocks to model context-dependent interactions in parallel. Finally, each sub-action’s representation is decoded separately, preserving tractability without assuming independence and maintaining permutation invariance.
We empirically evaluate SAINT on challenging benchmark tasks that exhibit both primarily state-independent and primarily state-dependent sub-action dependencies. Our results show that SAINT significantly outperforms strong baselines, including those with independent and autoregressive factorizations, across action spaces ranging from hundreds to over 16 million discrete actions. Targeted ablations validate the effectiveness of SAINT’s pre-attention FiLM-based state conditioning and quantify the computational overhead introduced by its attention mechanism. Together, these findings establish SAINT’s ability to efficiently learn coordinated behaviors in complex combinatorial action spaces.
@@ 2 Related Work
Combinatorial action spaces arise naturally in many sequential decision-making problems, including text-based interaction, games, and resource allocation. Prior approaches have introduced task-specific architectures or assumptions tailored to these settings [4, 5, 7, 13, 14, 29, 44]. Others exploit problem-specific structure, such as the amenability to gradually expanding curricula, to improve sample efficiency [10]. While effective in context, such methods typically lack generality and require significant manual design effort. In contrast, SAINT provides a domain-agnostic architecture that models sub-action dependencies through permutation-invariant self-attention, offering greater flexibility across structured decision problems. A parallel body of work addresses continuous control problems by discretizing the action space [2, 25, 36, 40], which contrasts with our focus on inherently discrete action spaces with combinatorial structure. Recent work in offline RL has explored learning in combinatorial action spaces under dataset constraints [3, 20, 35]; however, these approaches address challenges that differ from those in online learning.
A number of general-purpose architectures have been developed to scale RL to high-dimensional, combinatorial action spaces. One common strategy is to factorize the action space, treating each sub-action as conditionally independent given the state. This reduces the action space size from exponential to linear in the number of sub-actions [31, 38]. However, the independence assumption limits the expressiveness of the policy and breaks down in settings where sub-actions strongly interact. Autoregressive (AR) models [47] instead decompose the joint action distribution into a sequence of conditional distributions, allowing the policy to model dependencies among sub-actions. Yet, this introduces an arbitrary ordering over sub-actions — even when no natural order exists — violating permutation invariance and potentially impairing learning when the imposed structure misaligns with the task. The Wolpertinger architecture (Wol-DDPG) [8] embeds discrete actions into a continuous space and uses approximate nearest-neighbor search to retrieve a subset of candidate actions before value-based refinement. While scalable, this method performs poorly when the structure of the action space is inherently unordered [5], a limitation we verify experimentally.
Transformers in Reinforcement Learning
The success of the Transformer architecture [43] in sequence modeling has led to its adoption in RL, most notably through methods like Decision Transformer (DT) [6] and Trajectory Transformer (TT) [17]. These approaches use Transformer decoders to autoregressively predict actions based on past trajectories of states, actions, and returns, showing strong performance in offline RL.
Transformers have also been used in RL settings beyond trajectory modeling [1]. DTQN [9] applies a Transformer decoder for Q-learning in partially observable environments, while other work uses them for predictive world modeling [26, 34] and high-dimensional state representation learning [18, 19]. Attention mechanisms over state representations have also been shown to improve reasoning in complex environments. For example, multi-head attention in centralized critics enables agents to selectively aggregate information from others, improving stability and sample efficiency [16], while self-attention over learned entity embeddings supports relational inference and generalization [45].
In contrast to prior approaches that apply attention to model relations among external entities, agents, or trajectories, SAINT leverages self-attention over sub-actions within a single timestep. Inspired by the Transformer’s success in set processing [21], SAINT treats sub-actions as an unordered set and applies permutation-equivariant self-attention to model their interactions, enabling the policy to capture complex intra-action dependencies critical for combinatorial control.
@@ 3 Preliminaries
We consider RL problems formalized as a finite-horizon Markov Decision Process, defined by the tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},p,r,\gamma,\mu\rangle$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $p(s^{\prime}\mid s,a)$ the transition function, $r(s,a)$ the reward function, $\gamma\in[0,1]$ the discount factor, and $\mu$ the initial state distribution. A policy $\pi$ maps states to distributions over actions, $\pi:\mathcal{S}\rightarrow\mathbb{P}(\mathcal{A})$, thereby governing the agent’s behavior over episodes of length $H$.
In this work we assume actions have an explicit compositional structure. Specifically, the action space is a product of sub-action domains, $\mathcal{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{A}$, where each sub-action space $\mathcal{A}_{i}$ can be either discrete or continuous. A single action thus comprises $A$ sub-decisions, $\mathbf{a}=(a_{1},\dots,a_{A})$, with each $a_{i}\in\mathcal{A}_{i}$. This representation gives rise to high-dimensional action spaces with potentially complex dependencies among sub-actions.
@@
Attention is a general computational primitive that allows a model to selectively aggregate information from a set of inputs based on learned relevance scores [43]. Formally, given a set of queries $Q\in\mathbb{R}^{n_{q}\times d}$, keys $K\in\mathbb{R}^{n_{k}\times d}$, and values $V\in\mathbb{R}^{n_{k}\times d}$, the scaled dot-product attention computes output $\mathrm{Attn}(Q,K,V)=\text{softmax}(QK^{\top}/\sqrt{d})V$. Multi-head self-attention extends this by learning multiple independent projections and aggregating their outputs, enabling the model to capture different types of interactions in parallel. Transformers, which stack layers of multi-head self-attention with feedforward components, have been widely adopted in domains requiring flexible modeling of structured dependencies.
@@ 4 Sub-Action Interaction Network using Transformers (SAINT)
Figure 1: Overview of SAINT. Initial sub-action embeddings are conditioned on the state $s$ using FiLM. Stacked Transformer blocks then model dependencies among sub-actions. The resulting context-aware representations are passed to independent Decision MLPs, which output per-sub-action policy distributions used for factorized sampling.
Prior RL methods for learning in combinatorial action spaces typically represent sub-actions either as conditionally independent or in a fixed autoregressive sequence. These approaches limit policy expressiveness and are ill-suited for domains where sub-actions naturally form an unordered set, such as in traffic signal control. To address these limitations, we propose the Sub-Action Interaction Network using Transformers (SAINT), a novel policy architecture that models multi-component actions as unordered sets of $A$ sub-actions. SAINT first applies Feature-wise Linear Modulation (FiLM) [30] to inject global state information into sub-action embeddings drawn from a learned embedding table. These state-aware embeddings are then processed by multi-head Transformer blocks, enabling the model to capture higher-order dependencies among sub-actions. The resulting representations are passed through individual MLP decision networks to parameterize per-sub-action distributions, from which actions are sampled independently. This enables efficient action construction while preserving the ability to model complex interactions. To our knowledge, SAINT is the first policy architecture to model complex sub-action dependencies in combinatorial action spaces while preserving permutation equivariance. An overview of the SAINT architecture is shown in Figure 1.
@@
Given an environment state $\mathbf{s}\in\mathbb{R}^{d_{s}}$, SAINT begins by assigning each sub-action $i\in\{1,\dots,A\}$ a learnable embedding from a table $\mathrm{Embed}\in\mathbb{R}^{A\times d}$, yielding a base representation $\mathbf{e}_{i}=\mathrm{Embed}(i)\in\mathbb{R}^{d}$. Each sub-action index $i$ is thus mapped to a learnable $d$-dimensional vector, regardless of the original cardinality of $\mathcal{A}_{i}$, providing a shared embedding space that enables uniform processing by the subsequent Transformer layers.
@@
To condition the sub-action representations on the global state, we apply FiLM to the initial embeddings $\mathbf{e}_{i}$. A shared MLP $g:\mathbb{R}^{d_{s}}\to\mathbb{R}^{2d}$ maps the state $\mathbf{s}$ to FiLM parameters $(\boldsymbol{\gamma},\boldsymbol{\beta})=g(\mathbf{s})$, which are then used to apply a sub-action-wise affine transformation:
$\tilde{\mathbf{e}}_{i}=\boldsymbol{\gamma}\odot\mathbf{e}_{i}+\boldsymbol{\beta}\;.$
@@
Following FiLM-based conditioning, the matrix of state-aware sub-action representations, $\tilde{\mathbf{E}}\in\mathbb{R}^{A\times d}$, is processed by a stack of $L$ identical Transformer blocks, omitting positional encodings to preserve permutation equivariance. Letting $\mathbf{X}^{(0)}=\tilde{\mathbf{E}}$, each block $\ell=1,\dots,L$ first computes multi-head self-attention over the sub-actions. Specifically, it constructs queries, keys, and values via learned projections: $\mathbf{Q},\mathbf{K},\mathbf{V}=\mathbf{X}^{(\ell-1)}W^{Q},\mathbf{X}^{(\ell-1)}W^{K},\mathbf{X}^{(\ell-1)}W^{V}$, and applies scaled dot-product attention to model interactions among sub-actions. The output of the attention layer is passed through a position-wise feed-forward network (FFN) applied independently to each sub-action embedding. Each sub-layer, attention and FFN, is followed by residual connections and layer normalization, as in standard Transformer architectures. This mechanism enables the policy to model state-conditioned dependencies between sub-actions while maintaining permutation equivariance.
The final output $\mathbf{X}=\mathbf{X}^{(L)}\in\mathbb{R}^{A\times d}$ provides a representation $\mathbf{x}_{i}$ for each sub-action $i$ that integrates global state information from FiLM conditioning and inter-sub-action dependencies modeled via self-attention. A sub-action-specific decision MLP $f_{i}:\mathbb{R}^{d}\to\mathbb{R}^{P_{i}}$ then maps $\mathbf{x}_{i}$ to the parameters of a distribution $\pi_{i}$ over the sub-action space $\mathcal{A}_{i}$. For example, if $\mathcal{A}_{i}$ is discrete with $K_{i}$ possible actions, $f_{i}$ outputs $P_{i}=K_{i}$ logits for a categorical distribution $\pi_{i}(a_{i}\mid\mathbf{s})=\mathrm{Categorical}(f_{i}(\mathbf{x}_{i}))$. If $\mathcal{A}_{i}$ is continuous, $f_{i}$ may output $P_{i}=2$ values — the mean and variance of a Gaussian distribution.
Because each sub-action representation encodes state and interaction context, the policy can be factorized across sub-actions without loss of modeling capacity:
$\log\pi(\mathbf{a}\mid\mathbf{s})=\sum_{i=1}^{A}\log\pi_{i}(a_{i}\mid\mathbf{s})\;.$
This factorization is necessary in combinatorial action spaces, where evaluating or optimizing over the full joint distribution $P(a_{1},\dots,a_{A}\mid\mathbf{s})$ is typically infeasible.
A separate value network $V:\mathbb{R}^{d_{s}}\to\mathbb{R}$ provides a scalar estimate of the state value. In our implementation, $V$ is an MLP that conditions only on $\mathbf{s}$ and does not share parameters with the policy network, though shared encoders could be used for efficiency if desired. SAINT is agnostic to the choice of learning algorithm and can be integrated with standard policy gradient and actor-critic methods, such as PPO [33] and A2C [27], using their corresponding training objectives.
By combining FiLM-based state conditioning with permutation-equivariant self-attention, SAINT provides a unified architecture for learning policies in combinatorial action spaces. Its design enables rich, state-dependent interaction modeling among sub-actions while maintaining scalability through independent decoding. The resulting factorized policy is both expressive and tractable, capable of operating in high-dimensional action spaces without assuming independence or imposing arbitrary ordering. Together with a modular value function and compatibility with standard RL algorithms, SAINT offers a flexible and general-purpose framework for structured decision-making. Pseudocode for action selection using SAINT is provided in Algorithm 1.
Algorithm 1 Sub-Action Interaction Network using Transformers (SAINT)
1:Construct initial embedding matrix $\mathbf{E}\leftarrow[\text{Embed}(1),\dots,\text{Embed}(A)]^{\top}$triangleright $\mathbf{E}\in\mathbb{R}^{A\times d}$
2:Compute FiLM parameters from state $(\boldsymbol{\gamma},\boldsymbol{\beta})\leftarrow g(\mathbf{s})$
3:Apply FiLM conditioning element-wise $\tilde{\mathbf{E}}\leftarrow\boldsymbol{\gamma}\odot\mathbf{E}+\boldsymbol{\beta}$
4:Process through Transformer blocks $\mathbf{X}\leftarrow\text{TransformerBlocks}(\tilde{\mathbf{E}})$triangleright $\mathbf{X}\in\mathbb{R}^{A\times d}$
5:Initialize action vector $\mathbf{a}=(\dots)$
6:Initialize total log-probability $\ell\leftarrow 0$
7:for $i=1$ to $A$ do $\triangleright$ Independent decoding per sub-action
8:     Extract context-aware embedding for sub-action $\mathbf{x}_{i}\leftarrow\mathbf{X}[i,:]^{\top}$
9:     Compute distribution parameters using decision network $\text{params}_{i}\leftarrow f_{i}(\mathbf{x}_{i})$
10:     Define distribution $\pi_{i}\leftarrow\text{Distribution}(\mathcal{A}_{i},\text{params}_{i})$triangleright e.g., Categorical
11:     Sample action $a_{i}\sim\pi_{i}$
12:     Accumulate log-probability $\ell\leftarrow\ell+\text{log-prob}(a_{i};\pi_{i})$
13:     Store sampled sub-action $\mathbf{a}[i]\leftarrow a_{i}$
14:end for
15:return Return joint action $\mathbf{a}=(a_{1}..a_{A})$ and total log-prob $\ell$

### v2

# SAINT: Attention-Based Policies for Discrete Combinatorial Action Spaces

@@ Abstract
The combinatorial structure of many real-world action spaces leads to exponential growth in the number of possible actions, limiting the effectiveness of conventional reinforcement learning algorithms. Recent approaches for combinatorial action spaces impose factorized or sequential structures over sub-actions, failing to capture complex joint behavior. We introduce the Sub-Action Interaction Network using Transformers (SAINT), a novel policy architecture that represents multi-component actions as unordered sets and models their dependencies via self-attention conditioned on the global state. SAINT is permutation-invariant, sample-efficient, and compatible with standard policy optimization algorithms. In 20 distinct combinatorial environments across three task domains, including environments with nearly 17 million joint actions, SAINT consistently outperforms strong baselines.  [^1] Code is available at https://github.com/matthewlanders/SAINT
@@ 1 Introduction
Reinforcement learning (RL) has achieved remarkable success across a range of domains, primarily through methods designed for either small, discrete action spaces (Hessel et al., 2018; Mnih et al., 2015; van Hasselt et al., 2015; Mnih et al., 2016) or continuous control (Fujimoto et al., 2018; Haarnoja et al., 2018; Lillicrap et al., 2015; Schulman et al., 2017). Many real-world problems, however, involve action spaces that lie between these extremes. These large discrete combinatorial spaces are defined as Cartesian products of multiple subspaces, where each joint action $\mathbf{a}=(a_{1},\dots,a_{A})$ consists of several coordinated sub-actions. Such settings, which arise in critical applications like traffic signal control (Rasheed et al., 2020) and drug selection (Tang et al., 2022), require learning policies that can effectively represent and reason about exponentially large, structured action spaces.
Traditional RL methods model discrete action spaces with a flat categorical policy, but this becomes intractable in combinatorial settings where the number of actions scales as $\prod_{d=1}^{A}m_{d}$ for $A$ sub-action dimensions with $m_{d}$ choices each. To mitigate this combinatorial explosion, existing approaches (Dulac-Arnold et al., 2015; Pierrot et al., 2021; Tavakoli et al., 2018; Zhang et al., 2018) rely on simplifying assumptions that constrain the representational capacity of the policy class. One family of methods (Pierrot et al., 2021; Tavakoli et al., 2018) factorizes the policy as $\pi(\mathbf{a}\mid s)=\prod_{i}\pi_{i}(a_{i}\mid s)$, which cannot represent interactions between sub-actions. Another class of approaches (Zhang et al., 2018) imposes a fixed autoregressive order, specifying a policy class of distributions of the form $\pi(\mathbf{a}|s)=\prod_{i}\pi_{i}(a_{i}|s,a_{<i})$. This introduces an arbitrary sequence over sub-actions, breaking permutation invariance and impairing learning when the imposed order misaligns with the true dependency structure. Many real-world tasks violate both the independence and fixed-order assumptions. In healthcare, for example, drug combinations can exhibit complex interaction effects — treatments may be safe individually but harmful together, motivating permutation-invariant models for combinatorial action spaces. Our work targets precisely these settings: combinatorial action spaces wherein sub-action indexing is arbitrary or only weakly meaningful, and the fundamental structure lies in sub-action interactions rather than in any prescribed ordering.
We introduce the Sub-Action Interaction Network using Transformers (SAINT), a policy architecture that learns explicit representations of combinatorial actions by treating them as unordered sets of sub-actions. Through self-attention conditioned on the global state, SAINT captures dependencies among sub-actions to produce expressive yet tractable policies. SAINT proceeds in three stages. First, global state information is injected into initial sub-action representations. Next, self-attention (Vaswani et al., 2017) is applied over the set of state-conditioned representations to capture sub-action dependencies while preserving permutation equivariance. Finally, the representations are decoded in parallel, producing action distributions that preserve the modeled interactions while remaining computationally tractable.
We evaluate SAINT on challenging benchmark tasks, which exhibit both state-independent and state-dependent sub-action dependencies, including traffic light control (Zhang et al., 2019), navigation (Landers et al., 2024), and discretized MuJoCo locomotion tasks (Towers et al., 2024). Our results demonstrate that by learning a more expressive representation of the action space’s internal structure, SAINT consistently outperforms strong factorized and autoregressive baselines, scaling to environments with nearly 17 million discrete actions. Targeted ablations validate the role of state conditioning and show that the additional cost of modeling dependencies is often offset by substantial gains in sample efficiency. Together, these findings establish that learning explicit representations of sub-action interactions is a practical and scalable approach to decision-making in complex combinatorial domains.
@@ 2 Related Work
Combinatorial action spaces arise naturally in sequential decision problems such as traffic signal control, games, and resource allocation. Prior work has introduced task-specific architectures, imposed domain-specific assumptions, or exploited problem-specific structure (Bello et al., 2016; Chen et al., 2023; Delarue et al., 2020; He et al., 2015; 2016; Nazari et al., 2018; Zahavy et al., 2018; Farquhar et al., 2020). Such methods typically lack generality and require manual design effort. A parallel body of work addresses continuous control problems by discretizing the action space (Barth-Maron et al., 2018; Metz et al., 2017; Tang and Agrawal, 2020; Van de Wiele et al., 2020), which contrasts with our focus on inherently discrete action spaces with combinatorial structure.
A number of general-purpose architectures have been developed to scale RL to large combinatorial action spaces. One strategy reduces complexity by assuming conditional independence across sub-actions (Pierrot et al., 2021; Tavakoli et al., 2018), while another imposes an autoregressive order (Zhang et al., 2018). These approaches improve tractability but either ignore dependencies among sub-actions or introduce arbitrary orderings that break permutation symmetry. Retrieval-based methods such as Wolpertinger (Wol-DDPG) (Dulac-Arnold et al., 2015) scale to large discrete spaces by embedding and pruning candidate actions, but similarly fail to capture the joint structure of unordered sub-actions (Chen et al., 2023). These limitations motivate architectures that can represent dependencies across sub-actions while preserving permutation invariance.
Transformers for Action Representation
Efforts to use Transformers for action space modeling have largely focused on sequential representations. RT-1 (Brohan et al., 2022) and RT-2 (Zitkovich et al., 2023) tokenize robot control trajectories and decode action tokens autoregressively, while Q-Transformer (Chebotar et al., 2023) autoregresses across action dimensions within a timestep. Learned tokenizers such as FAST (Pertsch et al., 2025) compress high-frequency control signals into vocabularies for vision–language–action (VLA) training. Extensions of trajectory-based models (Chen et al., 2021; Shang et al., 2022), interleave state and action tokens or design state-aware tokenizations to handle multi-component actions. Together, these methods establish the efficacy of Transformers for action encoding, but have been mostly designed for continuous control tasks in offline RL or VLA settings that are not directly comparable to the combinatorial domains we study. Moreover, they largely rely on sequential decompositions that impose arbitrary order and obscure permutation symmetry.
@@ 3 Preliminaries
We consider RL problems formalized as a Markov Decision Process, defined by the tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},p,r,\gamma,\mu\rangle$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $p(s^{\prime}\mid s,a)$ the transition function, $r(s,a)$ the reward function, $\gamma\in[0,1]$ the discount factor, and $\mu$ the initial state distribution. A policy $\pi$ maps states to distributions over actions, $\pi:\mathcal{S}\rightarrow\mathbb{P}(\mathcal{A})$, which defines the agent’s behavior in the environment.
In this work we assume actions have an explicit compositional structure. Specifically, the action space is a product of sub-action domains, $\mathcal{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{A}$, where each sub-action space $\mathcal{A}_{i}$ is a discrete set. A single action thus comprises $A$ sub-decisions, $\mathbf{a}=(a_{1},\dots,a_{A})$, with each $a_{i}\in\mathcal{A}_{i}$. This representation gives rise to high-dimensional action spaces with potentially complex dependencies among sub-actions.
@@
Attention is a general computational primitive that allows a model to selectively aggregate information from a set of inputs based on learned relevance scores (Vaswani et al., 2017). Formally, given a set of queries $Q\in\mathbb{R}^{n_{q}\times d}$, keys $K\in\mathbb{R}^{n_{k}\times d}$, and values $V\in\mathbb{R}^{n_{k}\times d}$, the scaled dot-product attention computes output $\mathrm{Attn}(Q,K,V)=\text{softmax}(QK^{\top}/\sqrt{d})V$. Multi-head self-attention extends this by learning multiple independent projections and aggregating their outputs, enabling the model to capture different types of interactions in parallel. Transformers, which stack layers of multi-head self-attention with feedforward components, have been widely adopted in domains requiring flexible modeling of structured dependencies.
@@ 4 Sub-Action Interaction Network using Transformers (SAINT)
Figure 1: Overview of SAINT. Initial sub-action embeddings are conditioned on the global state $\mathbf{s}$ to produce state-aware representations. Stacked Transformer blocks then model dependencies among sub-actions. The resulting context-aware representations are passed to independent Decision MLPs, which output per-sub-action policy distributions used for factorized sampling.
We introduce the Sub-Action Interaction Network using Transformers (SAINT), a policy architecture that learns a state-conditioned, permutation-equivariant representation of the action space, enabling efficient computation of expressive action distributions. The SAINT architecture comprises three stages: (1) state conditioning, which injects global state information into sub-action representations; (2) interaction modeling, which applies self-attention to model higher-order relationships among sub-actions while preserving permutation equivariance; and (3) action decoding, which transforms each sub-action representation into a distribution over its discrete choices, with all sub-actions decoded in parallel to maintain tractability. An overview of the SAINT architecture is shown in Figure 1.
4.1 State Conditioning
@@
SAINT represents each sub-action $i\in\{1,\dots,A\}$ with a learnable embedding vector $\mathbf{e}_{i}=\mathrm{Embed}(i)\in\mathbb{R}^{d}$, drawn from a table $\mathrm{Embed}\in\mathbb{R}^{A\times d}$. With $d$ treated as a fixed hyperparameter shared across all sub-actions, each sub-action is represented by a $d$-dimensional embedding independent of its original cardinality $|\mathcal{A}_{i}|$, yielding a shared space for uniform processing by the subsequent Transformer layers.
Because sub-action identity alone is insufficient for decision-making, SAINT augments each base embedding $\mathbf{e}_{i}$ with information from the global state $\mathbf{s}\in\mathbb{R}^{d_{s}}$, enabling dependencies to be modeled in a state-aware manner. While several conditioning mechanisms such as cross-attention or concatenation are possible, we adopt Feature-wise Linear Modulation (FiLM) (Perez et al., 2018), which we found to be effective and parameter-efficient (see Appendix C). Notably, FiLM preserves the $d$-dimensional width of each sub-action embedding, introducing no additional projection dimensions. Its prior success in incorporating state information in RL (Brohan et al., 2022) further supports this choice.
@@
An MLP $g:\mathbb{R}^{d_{s}}\to\mathbb{R}^{2d}$ processes the global state $\mathbf{s}$ once to produce FiLM parameters $(\bm{\gamma},\bm{\beta})=g(\mathbf{s})$, which are then applied uniformly to all sub-action embeddings via an affine transformation:
$\tilde{\mathbf{e}}_{i}=\bm{\gamma}\odot\mathbf{e}_{i}+\bm{\beta}\;.$
4.2 Interaction Modeling
@@
The matrix of state-aware sub-action representations $\tilde{\mathbf{E}}\in\mathbb{R}^{A\times d}$ is then processed by a stack of $L$ Transformer blocks, with positional encodings omitted to preserve permutation equivariance. Letting $\mathbf{X}^{(0)}=\tilde{\mathbf{E}}$, each block $\ell=1,\dots,L$ performs multi-head self-attention followed by a feed-forward network (FFN). Specifically, queries, keys, and values are obtained by linear projections of the previous layer’s output:
$\mathbf{Q},\mathbf{K},\mathbf{V}=\mathbf{X}^{(\ell-1)}W^{Q},\;\mathbf{X}^{(\ell-1)}W^{K},\;\mathbf{X}^{(\ell-1)}W^{V}\;,$
which are then used in scaled dot-product attention to model interactions among sub-actions. The attention output is then passed through a position-wise FFN applied independently to each sub-action embedding. This design allows SAINT to model state-conditioned dependencies between sub-actions while maintaining permutation equivariance.
4.3 Action Decoding
In the final stage, each context-aware sub-action representation $\mathbf{x}_{i}$ is passed through a sub-action-specific decision MLP, $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K_{i}}$, which outputs a vector of $K_{i}$ logits. These logits are then transformed into a probability distribution over the $K_{i}$ discrete choices for sub-action $i$ via the softmax function. The resulting policy for sub-action $i$ is thus given by:
$\pi_{i}(a_{i}\mid\mathbf{s})=\mathrm{Categorical}(\text{softmax}(f_{i}(\mathbf{x}_{i})))\;.$
Because each sub-action representation $\mathbf{x}_{i}$ from the Interaction Modeling stage is conditioned on the global state and incorporates information from the other sub-actions, the policy can be expressed as independent sub-action distributions without loss of modeling capacity, preserving tractability in combinatorial spaces where representing the full joint distribution would be infeasible.
Algorithm 1 Policy Learning with the SAINT Architecture
1:Initialize SAINT policy $\pi_{\theta}$ and value network $V_{\phi}$
2:for each training iteration do
3:  Collect a batch of transitions $(\mathbf{s}_{t},\mathbf{a}_{t},r_{t},\text{done}_{t})$ by executing $\pi_{\theta}$
4:  Compute return $R_{t}$ and weighting term $w_{\Phi}(\mathbf{s}_{t},\mathbf{a}_{t},R_{t})$ for each transition
5:  Compute policy log-probabilities $\ell_{t}\leftarrow\textsc{LogProbs}(\mathbf{s}_{t},\mathbf{a}_{t})$ $\triangleright$ See function below
6:  Update policy $\theta$ by ascending the objective $\mathbb{E}_{t}[w_{\Phi}(\mathbf{s}_{t},\mathbf{a}_{t})\cdot\ell_{t}]$
7:  Update value function $\phi$ by descending the loss $\mathbb{E}_{t}[(V_{\phi}(\mathbf{s}_{t})-R_{t})^{2}]$
8:end for
9:
10:function LogProbs($\mathbf{S},\mathbf{A}_{\text{taken}}$)
11:  Get sub-action embeddings $\mathbf{E}\leftarrow[\text{Embed}(1),\dots,\text{Embed}(A)]^{\top}$
12:  Inject state information $\tilde{\mathbf{E}}\leftarrow\text{StateCondition}(\mathbf{S},\mathbf{E})$
[…]
14:  Get logits for each sub-action $\text{Logits}_{i}\leftarrow f_{i}(\mathbf{X}_{[:,i]})$ for $i=1,\dots,A$
15:  Compute log-probabilities $\log\mathbf{P}_{[:,i]}\leftarrow\log\pi_{i}(\mathbf{A}_{\text{taken}[:,i]}\mid\text{Logits}_{i})$ for $i=1,\dots,A$
16:  return $\sum_{i=1}^{A}\log\mathbf{P}_{[:,i]}$
17:end function
4.4 Compatibility with RL Algorithms
The SAINT architecture is compatible with any RL algorithm for which the actor objective maximizes the log-likelihood of sampled joint actions $\mathbf{a}$, weighted by some functional $w_{\Phi}(s,\mathbf{a})$:
$\max_{\theta}\;\mathbb{E}_{(s,\mathbf{a})\sim\mu}\big[w_{\Phi}(s,\mathbf{a})\log\pi_{\theta}(\mathbf{a}\mid s)\big],$
where $\mu$ denotes the sampling distribution over $(s,\mathbf{a})$, arising either from an online policy or from a fixed dataset in the offline setting. The weight $w_{\Phi}(s,\mathbf{a})\geq 0$ is an algorithm-dependent scalar, such as an advantage term or a score derived from $Q_{\Phi}$.
Compatible methods include standard online algorithms such as PPO (Schulman et al., 2017) and A2C (Mnih et al., 2016), as well as offline approaches such as IQL (Kostrikov et al., 2021) and AWAC (Nair et al., 2020). SAINT also supports selection-based actor updates as in BCQ (Fujimoto et al., 2019), where the policy is trained on candidate joint actions drawn from a dataset or proposal distribution. SAINT remains compatible even when $w_{\Phi}(s,\mathbf{a})$ is computed with a factorized critic, since the critic is used only to produce a scalar weight for each sampled joint action from $\mu$. The actor is always updated toward the observed joint action $\mathbf{a}$, not an action reconstructed or optimized over by the critic.
Incompatibility arises when the actor objective requires global operations over the entire combinatorial action space, such as $\mathbb{E}_{\mathbf{a}^{\prime}\sim\pi*\theta}[Q_{\Phi}(s,\mathbf{a}^{\prime})]$ or $\max_{\mathbf{a}^{\prime}}Q_{\Phi}(s,\mathbf{a}^{\prime})$. These operations are computationally intractable unless $Q_{\Phi}$ is factorized; however, this changes the structure of the actor target, decomposing it into uncoordinated per-dimension terms and discarding cross-dimensional structure. This breaks alignment with SAINT’s objective of modeling dependencies among sub-actions. SAINT’s learning procedure is provided in Algorithm 1.